Lakeitha earned $12,000 more in her accounting job this year because she received her CPA certification. She decided to invest the money, which was split between two different savings accounts. One account was a certificate of deposit that paid 7.25% annual interest. The other was a money market account that paid 5.4% annual interest. At the end of one year, she had made $730. How much did she invest in the money market account?
Answer by stanbon(48551)

Let amount invested at 7.25% be "x".
The amount invested at 5.4% will be 12000-x
Interest at 7.25% is 0.0725x
Interest at 5.4% is 0.054(12000-x)=648-0.054x
EQUATIOH:
Interest + Interest =730
0.0725x+648-0.054x=730
0.0185x=82
x=$4432.43 (amount invested at 7.25%)
12000-x=7567.57 (amount invested at 5.4%)
Cheers,
stan H.